OpenStudy (anonymous):

I' not really sure how to solve this one I would greatly appreciate the help [9.05] Solve 2x2 + 3x + 6 = 0. Round solutions to the nearest hundredth. x ≈ −2.64 and x ≈ 1.14 x ≈ −1.14 and x ≈ 2.64 No real solutions x ≈ −1.11 and x ≈ −4.89

OpenStudy (anonymous):

we can use the formula for qudratic equtions of the type ax^2+bx+c=0

OpenStudy (anonymous):

\[x=(-b \pm(\sqrt{b^2-4(a*c)})/2a\]

OpenStudy (anonymous):

this has no real solutions

OpenStudy (anonymous):

thank you but can you explain why please im not familiar with the formula

OpenStudy (anonymous):

ok..you saw that sqrt(b^2-4(a*c) part?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

we have b^2=3^2=9. 4(a*c) will be 4*2*6=48

OpenStudy (anonymous):

then the element inside sqrt.() will be negative...there is no root for a neg.number

OpenStudy (anonymous):

thank you very much i really appreciate the help

OpenStudy (anonymous):

you are welcome:D